### v1

# HyperSurf: Quadruped Robot Leg Capable of Surface Recognition with GRU and Real-to-Sim Transferring

Abstract
This paper introduces a system of data collection acceleration and real-to-sim transferring for surface recognition on a quadruped robot. The system features a mechanical single-leg setup capable of stepping on various easily interchangeable surfaces. Additionally, it incorporates a GRU-based Surface Recognition System, inspired by the system detailed in the DogSurf paper [1]. This setup facilitates the expansion of dataset collection for model training, enabling data acquisition from hard-to-reach surfaces in laboratory conditions. Furthermore, it opens avenues for transferring surface properties from reality to simulation, thereby allowing the training of optimal gaits for legged robots in simulation environments using a pre-prepared library of digital twins of surfaces. Moreover, enhancements have been made to the GRU-based Surface Recognition System, allowing for the integration of data from both the quadruped robot and the single-leg setup. The dataset and code have been made publicly available.
Keywords: Robotics, Quadruped Robot, Surface Recognition, Terrain classification, IMU, GRU
I Introduction
Quadruped robots get a lot of attention nowadays, due to their potential advantage over wheeled robots to navigate unstructured off-road terrains and urban environments where people live and work. This ability opens a lot of potential applications in different spheres like autonomous inspection, agriculture [2], delivery [3], assisting for visually impaired people, space exploration. However, performing some tasks demands perception of different terrains and consequent adeptness to them or awareness of the user about the type of a surface.
Employing quadruped robots as assistants requires an effective system, which is able to evaluate whether a traversed surface is dangerous or not in urban environments. One of the crucial criteria in this case is the robot’s ability to evaluate surface slipperiness. A slippery surface significantly increases the risk of severe injuries, e.g., hip and spine cord fractures from falls. The statistic estimations suggest that hip fracture rates will nearly double by 2050 compared to 2023 levels [4]. Consequently, an assistant quadruped robot must possess the advanced intelligence to warn people about slippery surfaces.
Figure 1: The experimental stand consists of a leg, fixing stand, control electronics, IMU sensor, and the surface sample under investigation.
Investigating the surfaces of celestial bodies involves traversing regolith, which is the predominant material covering the surfaces of Mars, the Moon, and various other celestial bodies. The regolith is often loose, granular, and heterogeneous, posing challenges for traversal with wheeled robots. [5]. Therefore, using quadruped robots to explore the complex surfaces of celestial bodies is rational due to their ability to function and traverse intricate landscapes. However, locomotion in complex environment demands from legged robots to adapt their behavior to different parameters of surfaces, which stimulate to develop different approaches for surface recognition.
Our method builds upon the DogSurf [1] surface recognition system and expands the effectiveness of its method. In this paper, we have shifted our approach to dataset collection from gathering data by walking on a quadruped robot to collecting data in a laboratory setting while the robot stands on a single leg. This method offers advantages in terms of speed, simplicity, and cost-effectiveness, enabling faster data collection and experimentation, because our new one-leg setup does not require investment of manpower and time for datasets collections. Our GRU-based system offers a cost-effective solution that can be easily implemented across a wide range of legged robotics. Utilizing this setup would facilitate the collection of datasets, which could then be transferred to a quadruped robot or in simulation environment allowing to develop optimal locomotion patterns in simulation. By incorporating feedback from IMU sensor systems and adjusting parameters, this approach can enhance the adaptability of legged robots.
II Related Works
DogSurf is a novel method employing quadruped robots to aid visually impaired individuals in real-world navigation. This approach empowers the quadruped robot to identify slippery surfaces and provide auditory and tactile cues to signal the user when to halt. Additionally, a cutting-edge GRU-based neural network architecture boasting a mean accuracy of 99.925% was introduced for the multiclass surface classification task tailored specifically for quadruped robots. The dataset used for training was gathered using a Unitree Go1 Edu robot.
In addition to DogSurf, several other methodologies for surface recognition are available or currently employed in quadruped robot locomotion.
First approach involves employing various tactile estimation methods. Weerakkodi et al. [6] introduced the utilization of Touch Sensitive Foot (TSF) tactile sensors, positioned on the dog’s feet to gather data, subsequently processed by Convolutional Neural Networks (CNN). This method enables the attainment of a validation accuracy of 74.37% and a peak recognition score of 90% for line patterns. Despite its potential, this approach faces several limitations, such as the high service cost due to direct sensor contact with the surface, leading to gradual wear and tear over time.
Second approach involves utilizing audio data. Dimiccoli et al. [7] employed a gripper to manipulate objects and captured audio signals, subsequently utilized for CNN training, resulting in approximately 85% accuracy. Another successful attempt was made by Vangen et al. [8], who suggested employing a sensorized paw for audio-based terrain classification, achieving around 78% accuracy. However, this method exhibits limited real-world applicability due to the diverse ambient sounds present, which notably degrade the neural network’s accuracy.
Third approach involves utilizing Force/Torque (F/T) sensor data either independently or in conjunction with the aforementioned methods. Bednarek et al. [9] advocated for the utilization of data solely from F/T sensors on the quadrupedal robot ANYmal [10]. Additionally, they introduced a novel CNN-1d convolution architecture and devised a clustering method for terrain classification, achieving an accuracy of 93%. Another method employing a Transformer architecture was proposed in [11] and attained an accuracy of 97.33% on a QCAT dataset. Kolvenbach et al. [12] recommended combining F/T and IMU sensors for the ANYmal robot and developed a specialized test stand, achieving an accuracy of 98.6%.
On average, the IMU-based approach yields superior results and offers advantages such as no direct environmental contact, resilience to weather and time-of-day variations, and cost-effectiveness. This method has been successfully implemented in a walking robot utilizing gated recurrent units (GRU) [13].
III System Overview
III-A Hardware Architecture
For conducting experiments to determine surface types, a special experimental setup was created to ensure repeatability and uniformity of the data obtained when interacting with various surface types. The creation of the stand significantly simplified and reduced the cost of the experiment, as it eliminated the need to use a full-fledged quadruped robot. The main components of the stand (1) include: a full-size leg of a quadruped robot with three degrees of freedom, mounted on a vertical stand, control electronics, an IMU sensor, and the surface for investigation.
Figure 2: HyperSurf system overview, including modules and interfaces developed for surface recognition.
The leg mounting system allows it to move freely along the vertical guides, ensuring stability and repeatability of movements. Following control commands from the electronic unit, the leg joints are simultaneously bended and extended, allowing the leg to perform jumps in place to a height of about 20 mm, pushing off from the surface under investigation. At the end of the leg, an IMU sensor (HWT905-RS485) is attached to collect feedback.
The leg was designed and manufactured as part of a project to create an autonomous quadruped robotic system. It consists of 3 movable joints and 3 BLDC electric motors that drive its movement. The joint in contact with the surface is equipped with a resilient silicone tip to cushion the contact.
The control of electric motors is performed using an algorithm in the micro-ROS environment on the Intel NUC microcomputer. This same computer is responsible for collecting data from the IMU sensor. The entire system is designed for reliable and stable reproduction of the jumping pattern specified by the user on the surface under investigation.
III-B HyperSurf Control System
In the control system of the HyperSurf robotic leg, a combination of ROS2 and micro-ROS technologies governing its movements. Within the ROS2 package, specific leg trajectories are defined within Python scrips encapsulated in ROS2 packages. These trajectories, created to facilitate leg movement, are translated into motor angles, shaping the physical actions of the robotic limb. The computational required for these calculations is executed on an Intel NUC computer.
Once the trajectory calculations are finalized, they are converted into motor angles by an inverse kinematics algorithm. These angles, along with velocity and the coefficients $K_{p}$ desired position stiffness and $K_{d}$ velocity gain, are transmitted to the leg actuators via the /Hyper dog joints/Commands topic. The stiffness and damping coefficients were selected in advance, taking into account the load on each motor. They determine how quickly the desired position will be achieved and how rigidly this position will be held when external forces are applied. Sending data to the motors is feasible due to the created custom ROS2 message. The data is transmitted in three separate messages due to the leg’s configuration with three motors, each motor possessing a unique identifier (can_id). This enables the distinction between signals sent over a single CAN bus.
Micro-ROS agent is facilitating the seamless interaction between the high-level computation environment and the low-level hardware control. This agent establishes a vital link between the ROS2 ecosystem running on the computer and the microcontroller system, the /joints/Commands topic published by the ROS2 environment is subscribed to by the /STM_node, which is generating PWM signals for BLDC motors drivers.
HyperSurf not only executes the prescribed leg trajectories but also interfaces with motor encoders, capturing critical feedback on the limb’s position and velocity. This encoder data is meticulously processed, providing insights into the leg’s current state and ensuring precise control over its motion.
III-C Choice of Surfaces for Investigation
To reliably verify the idea of surface recognition based on feedback from the IMU sensor, a diverse set of surfaces with varying properties were assembled. The interaction of the robot leg was analyzed on four types of surfaces: a sheet of EVA polymer (A), a textured ceramic surface (B), a smooth enamel-coated ceramic tile surface (C), and carpet (D).
III-D Dataset Structure
Throughout the study, a dataset comprising over 1,000,000 samples (equivalent to approximately 1,5 hours of data) was amassed. Data collection encompassed four classes, including a sheet of EVA polymer, a textured ceramic surface, a smooth enamel-coated ceramic tile surface, and carpet. During the collection process, the stand operated in automatic stepping mode, varying its frequency to ensure an adequate diversity of data.
III-E GRU-based Surface Recognition
Our previous method DogSurf module utilizes a GRU neural network as the foundation. The IMU data from the accelerometer and gyroscope, is collected and appended to a queue containing 100 samples. During each iteration, a Forward Pass is conducted through the firstly Standart Scaler, then through Principal component analysis (PCA) transform, and then trough bidirectional GRU - neural network.
Figure 3: Neural network architecture
Code can be accessed by link: https://github.com/Eterwait/DogSurf. The results of DogSurf module performance of the HyperSurf compared to the State-of-the-Art are listed in I.
TABLE I: Comparison of HyperSurf with the State-of-the-art Approaches
Model Accuracy
Tactile Weerakkodi et al. [6] 0.74370
Audio Dimiccoli et al. [7] 0.84700
Vangen et al. [8] 0.77900
F/T HAPTR2 [11] 0.97330
Jakub Bednarek et al. [9] 0.93590
F/T & IMU Kolvenbach et al. [12] 0.98600
IMU Lomio et al. [14] 0.64950
Singh et al. [15] 0.88000
DogSurf 0.99925
For HyperSerf we get the new results which represented in Table II and confusion matrix in Fig. 4.
Figure 4: Confusion matrix for HyperSurf
TABLE II: Results for different surfaces for HyperSurf
Precision Recall F1 score
Carpet 0.99467 0.98902 0.99183
Rubber 0.98032 0.99032 0.98529
Tile 0.99695 0.99506 0.99600
Rough tile 1.00000 1.00000 1.00000
The outstanding results of surface recognition open new perspectives of utilizing quadruped robots both as companions, and also in complex landscapes.
IV Creating a Digital Twin of HyperSurf System
To conduct the stand simulation in the Gazebo environment, a digital twin was created based on the following steps:
IV-A Preparation of the stand description file
The three-dimensional model of the stand was meticulously crafted within Computer-Aided Design (CAD) software, leveraging a simplified rendition of the complete quadruped robot designed in Siemens NX. This comprehensive model encompasses all integral components essential for the stand’s functionality.
To seamlessly integrate the CAD model into the Gazebo simulator, a Unified Robot Description Format (URDF) was meticulously generated. Crafted through Fusion 360, this URDF file intricately delineates all stand components, their interactions, and kinematic properties. Notably, the URDF generation process was facilitated by a Python script available at the repository [16].
In line with the real-world stand setup, an IMU sensor was seamlessly incorporated into the dog’s leg using ROS tools. This addition faithfully mirrors the physical stand’s configuration, thereby enhancing the fidelity and realism of the simulation environment.
Furthermore, to simulate various surfaces within the Gazebo environment, including tiles, rubber, and brick surfaces, meticulous specifications were defined. These surface characteristics were meticulously outlined in the URDF file, with parameters carefully tuned to replicate physical attributes. Notably, parameters such as the friction coefficient, akin to a silicone surface, damping coefficient, and stiffness coefficient, were precisely calibrated to ensure accurate simulation dynamics.
IV-B Kinematics Control Implementation
Leg kinematics control was realized using symbolic and numerical computation techniques. Forward kinematics determined leg component positions, while inverse kinematics adjusted joint angles for desired end-effector positions. Symbolic computation with SymPy calculated transformation matrices and coordinates, while CasADi handled numerical optimization for joint angle adjustments. This approach enabled precise leg movement control, facilitating effective interaction with the environment and data collection from various surfaces within the Gazebo simulator.
IV-C Dataset Collection
In the Gazebo simulator, utilizing the digital twin of the stand, a dataset was gathered to train the surface recognition model. Similar to the real stand, data from the IMU sensor was collected and analyzed. This process aimed to demonstrate the transfer ability of data between the physical and simulated environments, showcasing the bidirectional capability of real-to-simulation and simulation-to-real data integration.
Figure 5: Digital twins of the one-leg stand in Gazebo simulator.
Figure 6: Comparison of the foot angular velocities during the jumps of simulated model and the real robotic leg.
V Novelty of HyperSurf System
V-A Boost of datasets collection
In comparison to collecting datasets on a quadruped robot, utilizing a one-leg test setup could increase the speed of GRU analysis by 7.5 times on average. It was achieved because of the rapid system deployment and autonomous operation in laboratory conditions.
V-B Shortening the Team for Collecting Datasets
Reducing the required collection time and utilizing a static location allowed for fewer people to be involved in the experiments. Whereas in previous work, more than five people were required due to the lengthy data collection process, in this study, only two researchers were sufficient to collect the necessary amount of data.
V-C Transferring Data
Utilizing data obtained from an IMU sensor on the stand, empirical modeling of a corresponding surface was conducted within the Gazebo simulation environment, as depicted in Figure 5. The surface modeling results within the Gazebo simulator are illustrated in Figure 6. When comparing the angular velocity of the leg between the data collected from the stand and the simulation, a simulation data accuracy of 98% was achieved.
The GRU model trained on the stand demonstrated seamless applicability to its digital twin within the simulation environment, and conversely, without necessitating additional preparation, achieving remarkable precision.
This provides ample opportunities for utilizing our system on various robotic platforms and enables the preparation of models for surfaces for which collecting datasets is challenging.
VI Conclusion and Future Work
A HyperSurf leg test setup has been developed, significantly enhancing surface recognition by assisting the DogSurf classification model, achieving a precision of slippery surface recognition up to 99.9%. It enables a reduction in the time required for data collection and minimizes the need for human involvement. Additionally, it facilitates data preparation under laboratory conditions. Furthermore, we have demonstrated that the prepared data can be easily transferred to simulations, which could greatly contribute to the utilization of this method in quadruped robotics.
In the future we aim to demonstrate how our HyperSurf surface recognition system can enhance the functionality of quadruped platforms. By utilizing feedback from the GRU and process it with DogSurf architecture, we can adapt gaits to different surfaces, adjust parameters of BLDC motors, and incorporate additional mechanics to improve both the energy consumption of the robot and its capability to traverse various terrains. Furthermore, we plan to continue developing our quadruped platform (Fig. 7) and integrate our system into it, along with all features based on surface recognition feedback. Additionally, we intend to extend our system to the area of object grasping, as leveraging information about surface properties of an object could significantly enhance the effectiveness of grasping.
Figure 7: 3D model of the HyperSurf quadruped robot.